OpenStudy (anonymous):

So I've tried looking for a lot of examples online and couldn't find any to get me jump started on a practice problem that I've no idea to do. I'm suppose to find the focus,vertex, directrix and axis of symmetry of the parabola with the equation Y = x^2 + 3 but no idea how if anyone could actually walk me through the steps so I can do others I would be very thankful.

Directrix (directrix):

@Lyo I think it would help to look at the relationship of these "parabola parts" which you are asked to find. I'll attach a diagram. Here is the link to the diagram if you want to see it in context. http://hotmath.com/hotmath_help/topics/directrix.html

Directrix (directrix):

So, you see what we are asked to find. Tell me, do you have a general equation of a parabola that you have been using in your class or one that you are supposed to use with this problem?

OpenStudy (anonymous):

Thank you, I am just starting on parabolas, I am not sure if it's what you're asking for but I was given two equations: (y - k)^2 = 4p(x - h) and (x -h )^2 = 4p(y - k) I am currently reading over the link you sent me. The problem doesn't require a specific equation as long as I can show my work and provide accurate answers.

Directrix (directrix):

Your given equation y = x^2 + 3 has the x term squared so use the general form that has a linear y and a squared x term. I have posted it below. I turned it around (symmetric property) to the way I like to look at it. I also divided both sides by (4p). I showed the steps below. (x -h )^2 = 4p(y - k) 4p(y - k) = (x -h )^2 (y - k) = (1/(4p) ) * (x -h )^2

Directrix (directrix):

So, we are starting with this: (y - k) = (1/(4p) ) * (x -h )^2 and comparing y = x^2 + 3 with it.

Directrix (directrix):

(y - k) = (1/(4p) ) * (x -h )^2 y - 3 = 1 * (x - 0)^2 (h,k) is the vertex so the vertex of the equation you were given is (0,3). @Lyo Do you agree with the vertex coordinates? Actually, we did not need this bulky formula to know that.

OpenStudy (anonymous):

From what I understand, Yes I agree but I do have a question what are we putting in for 1/4p to get 1, I understand everything else being plugged in just not that.

Directrix (directrix):

We are doing all that but because you posted that you had been reading examples and still were not quite sure what to do, I thought we would do this problem one step at a time. So, we will be getting to the value of p. But, without the vertex, the value of p will not help with the directrix and focus.

Directrix (directrix):

One more comment about the vertex. We could have taken this: Y = x^2 + 3 and noticed that the vertex had to be at (0,3) because the smallest x can be is 0 and then the corresponding y is 3. So, the vertex is a minimum point on the graph and the parabola opens upward.

Directrix (directrix):

Going back to this: (y - k) = (1/(4p) ) * (x -h )^2 y - 3 = 1 * (x - 0)^2 Notice that I placed a "one" in front of the (x-0)^2. That does not change the given equation.

Directrix (directrix):

1/(4p) = 1 So, 4p = 1 p = 1/4 @Lyo It is very important that you understand how the value of p was computed. So, look at it and see if you agree that p = 1/4 or if you have questions.

OpenStudy (anonymous):

I understand it, to get P we have to isolate it from everything else right?

Directrix (directrix):

Yes, after getting the given equation in the form so that we know the coefficient of the squared x component. That just happened to 1 this time.

Directrix (directrix):

Recall the definition of a parabola: A parabola is set of all points in a plane which are an equal distance away from a given point and given line. The point is called the focus of the parabola, and the line is called the directrix.

Directrix (directrix):

The vertex is a point on the parabola. So, it is equidistant from the directrix and the focus.

Directrix (directrix):

The focus and the vertex lie on the same vertical line. In this case, that line is x = 0. From the vertex (0,3) move up 1/4 (.25) units to get the coordinates of the focus to be (0,3.25).